OpenStudy (anonymous):

Graph using the Slope and the Y-Intercept. 4x-y=-4

OpenStudy (anonymous):

Help

OpenStudy (dumbcow):

put it in slope-intercept form y = mx + b

OpenStudy (dumbcow):

m=slope b = y-intercept

OpenStudy (dumbcow):

do you know how to solve equation for 1 variable? in other words get y by itself

OpenStudy (anonymous):

No, I really need help. How would I put the numbers in the formula?

OpenStudy (dumbcow):

theres no formula, just an equation you just have to move things around so it looks like y=mx +b its a 2-step process 1) move x term to other side, flip the sign 2) divide by coefficient of y on both sides, include every term

OpenStudy (anonymous):

thansk

OpenStudy (dumbcow):

your welcome what did you get y = ?

OpenStudy (anonymous):

im still working on it, but am a little confused, i need to find -4x-y=-4 but i dont know where to Put the numbers in the formula? What did you get? So this way i can go over mines and fix the errors that i have comitted,

OpenStudy (dumbcow):

ok the goal is to graph this line, 4x -y=-4 represents a line so we are not plugging in any numbers into a formula use algebra to get y by itself, then you can identify slope, y-intercept this then allows you to plot the points needed to graph the line

OpenStudy (dumbcow):

if you have to, just make a x,y chart and plot points that way to get your line

OpenStudy (anonymous):

oh okay

OpenStudy (dumbcow):

if x = 1, what does y equal ?

OpenStudy (anonymous):

-8?

OpenStudy (anonymous):

rite?

OpenStudy (dumbcow):

i get positive 8 4(1) -y = -4 -y = -8 y = 8

OpenStudy (anonymous):

oh okay, im sorry i forgot about the negative signs.

OpenStudy (dumbcow):

pay attention to those neg signs ok now pretend the 1 is x what does y equal

OpenStudy (dumbcow):

4x - y = -4 subtract 4x on both sides -y = -4x -4 divide by -1 y = 4x + 4 this is slope intercept form m = 4 b = 4 Now you can graph it